# Simulating Longitudinal Data from Marginal Structural Models

### v1

@@ 6.1 Simulation algorithm
We must then obtain $F(Y_{k+1}\mid\overline{L}_{j},do(\bar{A}_{k}),Y_{k}=0)$ for each $j$ to use in the next iteration. This can be done using (21), and hence, we can move to the $(k+1)$th iteration. By the principle of induction, the statement is true for all $k\geq 0$. ∎
In this basic setting, simulating $K$ time steps for each individual requires $\sum_{k=0}^{K-1}(2k+4)=K^{2}+3K$ calculations. This implies that the computational complexity grows quadratically with $K$. If an individual ceases to survive at any time step, the simulation is terminated for them.
Building upon this basic case, we analyze various extensions:
  • Multiple time-varying confounders: If the number of time-vary confounders increases, making $\bm{L}$ a vector of $p$ variables, the number of calculations for the distributions of ¯$\bm{\bar{}}L_{k-1}$, and $F(Y_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0)$ among the survivors at $k$ scales proportionally with $p$. The total number of computations becomes $\sum_{k=0}^{K-1}(2pk+3p+2)=pK^{2}+(2p+2)K$. Hence the computational complexity grows linearly with $p$. In the demonstration presented in the next subsection, we simulate data where there are two time-varying confounders.
  • Multiple outcomes: The approach to simulating additional outcomes depends on whether they are are survival-related or non-survival-related.
    • For non-survival outcomes, such as weight loss, simulated from marginal structural models – as opposed to from conditional distributions) – the number of calculations to derive $F(\bm{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y^{\triangle}_{k}=0)$ increases proportionally with the number of outcomes, $m$. Here, $\bm{Y}_{k}$ is a m-dimensional vector, and $Y_{k}^{\triangle}$ denotes the primary survival outcome. Specifically, the total number of calculations becomes $\sum_{k=0}^{K-1}(pmk+pk+pm+p+m+1)=\frac{mp+p}{2}K^{2}+(mp+m+p+1)K$.
    • For survival outcomes, such as competing risks (e.g., death from other causes), a hierarchical simulation approach is adopted. This involves defining a hypothetical sequence of events, where any order can be chosen, though some may be more intuitive than others. For instance, the simulation first evaluates death from the primary cause; if an individual survives, it then assesses death from other causes. The distribution of ¯$\bm{\bar{}}L_{k-1}$ is calculated sequentially conditional on surviving from both risks according to this order. As a result, the number of calculations scales with $m$. Additionally, similar to the non-survival case, deriving $(F(\bm{Y}_{k+1}\mid do(\bar{A}_{k}),\bm{Y}_{k}=0)$ requires $m$–fold more computations. In total, $\sum_{k=0}^{K-1}(pmk+pmk+pm+p+m+1)=mpK^{2}+(mp+m+p+1)K$ calculations are needed.
    In both scenarios, the computational complexity grows linearly with the number of outcomes, $m$.
  • Censoring: It is customary to specify and simulate censoring $M_{k}$ using a full conditional distribution $F(M_{k+1}\,|\,\bar{L}_{k},do(\bar{A}_{k}),\bm{Y}_{k}=0)$, which represents the likelihood of an individual being censored given their past history. This requires adding one additional calculation at the end of each time step to simulate $M_{k+1}$.
These calculations are all evaluations of $h$-functions and their inverses. For commonly used copulas with analytic $h$-functions, these evaluations are generally efficient and computationally fast.
6.2 Demonstration
We use an example to demonstrate the proposed algorithm simulating longitudinal time-to-event data. This setup broadly follows the simulation study in Seaman and Keogh (2024).
We consider baseline variables $B$, $C$ and two time-varying covariates $L$ and $W$. $A_{k}$ indicates treatment at follow-up $k$. Let the time to event be $T$ and $Y_{k}=I(T>k)$ be the survival status observed at each visit. Specifically we choose that at baseline:
@@ 6.2 Demonstration
$C\sim\operatorname{Bernoulli}(0.7)$ $W_{0}\sim\operatorname{Bernoulli}(\operatorname{expit}\{-0.2+0.5\,B+0.5\,C\})$
@@
We assume a Cox marginal structural model given by:
$\lambda(t\,|\,do(\bar{A}_{\left\lfloor t\right\rfloor}),C)=\lambda_{0}\exp\left(\beta_{C}\,C+\beta_{A}\,A_{\left\lfloor t\right\rfloor}+\beta_{CA}\,C\,A_{\left\lfloor t\right\rfloor}\right),$ (25)
where $\lambda_{0}$ is the baseline hazard and is assumed to be constant over time, $A_{\left\lfloor t\right\rfloor}$ represent the treatment status at the most recent visit prior to time $t$ (with $\left\lfloor t\right\rfloor$ denoting the largest integer less than $t$). We set $\lambda_{0}=\exp(-2)$ and $\beta_{C}=-0.5$. We vary the values of $(\beta_{A},\beta_{CA})$ to construct scenarios with different strengths of treatment: a strong effect scenario $(-0.5,0.3)$ and a weak effect scenario $(0.2,0.1)$. This parameterization assumes a constant hazard rate between follow-up visits, influenced by the treatment status at the most recent visit. This implies that given survival at follow-up time $k$, the survival time $T$ follows an exponential distribution with the rate equal to the hazard above:
$T-k\,|\,T>k\sim\operatorname{Exp}(\lambda(t\,|\,do(\bar{A}_{\left\lfloor t\right\rfloor}),C)).$
If $T-k<1$ then $Y_{k+1}=1$, and 00 otherwise. We set the total number of follow-ups $K$ to 6. For simplicity we set the copulas that weave the dependence between $L$, $W$ and $Y$ so those that describe the dependency between $L_{j}$ and $Y_{k+1}$ for $j=0,1,2,\dots,k$ are Student t with parameter $\rho_{L}$ and 5 degrees of freedom, and those for $W_{j}$ and $Y_{k+1}$ (given $L_{k}$) for $j=0,1,2,\dots,k$ are Gaussian with correlation $\rho_{W}$. We vary the values of $\rho_{L}$ and $\rho_{W}$ to construct scenarios with different strengths of confounding. Table 4 summarizes the different parameter values. In total, we consider 6 scenarios corresponding to the combination of different effect size and confounding scenarios, and three sample sizes: $n=\{1000,2000,5000\}$.
@@
Remark 6.2.
The proposed algorithm, as previously discussed, exclusively relies on evaluating analytic functions, making it highly efficient. Table 5 reports the runtime required to simulate data under the setup for the large causal effect with medium confounding scenario. The simulations were carried out on a local laptop using a single score without parallelization. Generating data for $n$ up to 10,000 units takes only a few seconds, while even for a very large sample size ($n=50,\!000$) and 11 time steps ($K=10$), the runtime remains under one minute. This computational efficiency is crucial for tasks such as method evaluation and sample size calculations in clinical trial design, where simulations must be repeated across various parameter values. For example, determining the sample size might require testing 10 different parameter combinations (such as effect sizes, prevalence of drop-in treatments, or adherence levels) across four sample sizes, each bootstrapped 200 times to estimate statistical power. This requires $10\times 4\times 200=8,\!000$ simulations. Using our method, suppose $n=10,000$ and $K=10$, the total runtime is approximately 13 hours on a single core, which can be reduced to about 2 hours using an 8-core laptop or under 10 minutes on a 50-core CPU through parallelization.
@@
We contrast the results of two approaches for fitting the Cox marginal structural model specified in (25). The first employs stabilized inverse probability of treatment weighting, with weights for subject $i$ at time step $k$ calculated as:
@@
The other method is to fit model (25) nominally, without weights or any adjustment for confounding. Table 6 summarizes the bias of estimates produced by inverse probability of treatment weighted and unweighted Cox regressions, from 500 replications. The biases are compared with their corresponding Monte Carlo standard errors as a measure of uncertainty. As the sample size increases, generally, the biases of estimates for $\beta_{C}$, $\beta_{A}$ and $\beta_{AC}$ obtained with inverse probability weights converge to zero, with their absolute values substantially smaller than the corresponding Monte Carlo standard deviations. In contrast, the unweighted estimates exhibit considerable bias, particularly for the estimate of the treatment effect, $\beta_{A}$, where the bias does not diminish as $n$ increases to $5000$. The magnitude of this bias far exceeds the Monte Carlo standard errors. The unweighted estimates are most biased under the high level of confounding by $L$ and $W$, since this is not adjusted for.
Level of confounding Effect size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ IPTW Unweighted IPTW Unweighted IPTW Unweighted High $1000$ $-0.04(2.27)$ $0.27(3.09)$ $3.73(1.73)$ $-1.05(1.54)$ $0.49(2.22)$ $2.31(9.34)$ $\beta_{A}=-0.5$ $2000$ $-0.07(1.68)$ $0.02(2.09)$ $3.81(1.21)$ $-1.03(1.06)$ $0.34(1.51)$ $1.63(9.52)$ $5000$ $0.02(1.25)$ $-0.09(1.58)$ $3.89(0.76)$ $-0.96(0.68)$ $0.24(0.97)$ $1.08(9.44)$ $1000$ $-0.03(2.35)$ $0.03(2.78)$ $3.64(1.46)$ $-0.93(1.59)$ $0.44(1.94)$ $1.99(9.62)$ $\beta_{A}=0.2$ $2000$ $-0.11(1.60)$ $0.17(2.02)$ $3.65(1.18)$ $-0.93(1.11)$ $0.49(1.51)$ $1.49(9.46)$ $5000$ $-0.03(1.14)$ $0.07(1.36)$ $3.65(0.71)$ $-0.95(0.71)$ $0.51(0.89)$ $1.01(9.38)$ Medium $1000$ $-0.07(2.04)$ $0.23(2.91)$ $1.90(1.73)$ $-0.43(1.51)$ $0.08(2.22)$ $2.19(9.48)$ $\beta_{A}=-0.5$ $2000$ $-0.02(1.60)$ $0.09(2.27)$ $1.95(1.22)$ $-0.35(1.01)$ $-0.01(1.55)$ $1.65(9.52)$ $5000$ $-0.09(1.10)$ $0.07(1.59)$ $1.97(0.80)$ $-0.42(0.64)$ $-0.02(1.03)$ $1.21(9.32)$ $1000$ $-0.12(2.06)$ $0.14(2.70)$ $1.75(1.54)$ $-0.36(1.43)$ $0.09(1.92)$ $2.02(9.50)$ $\beta_{A}=0.2$ $2000$ $-0.10(1.62)$ $0.09(2.09)$ $1.68(1.11)$ $-0.42(1.08)$ $0.17(1.44)$ $1.49(9.36)$ $5000$ $-0.03(1.04)$ $0.03(1.32)$ $1.78(0.68)$ $-0.39(0.66)$ $0.14(0.88)$ $0.95(9.54)$ Low $1000$ $0.01(1.82)$ $0.21(2.90)$ $-1.95(1.83)$ $0.46(1.29)$ $-0.20(2.20)$ $2.36(9.50)$ $\beta_{A}=-0.5$ $2000$ $-0.10(1.32)$ $0.07(2.13)$ $-1.83(1.25)$ $0.44(0.89)$ $-0.30(1.59)$ $1.69(9.58)$ $5000$ $-0.02(1.04)$ $0.03(1.46)$ $-1.77(0.84)$ $0.47(0.58)$ $-0.32(0.99)$ $1.07(9.42)$ $1000$ $0.04(1.98)$ $-0.03(2.65)$ $-1.88(1.46)$ $0.53(1.30)$ $-0.31(1.88)$ $1.98(9.64)$ $\beta_{A}=0.2$ $2000$ $-0.02(1.45)$ $0.01(1.87)$ $-1.85(1.10)$ $0.43(0.99)$ $-0.26(1.38)$ $1.43(9.44)$ $5000$ $-0.01(0.92)$ $-0.06(1.25)$ $-1.89(0.68)$ $0.43(0.59)$ $-0.25(0.84)$ $0.95(9.50)$
Table 7 presents the coverage rates of the 95% confidence intervals for both the inverse probability weighted and unweighted estimators. Consistent with the illustrative example in Section 5, the confidence intervals for the inverse probability weighted estimator were computed using two methods:one based on sandwich variance estimation and the other on bootstrap resampling. The results indicate that the confidence intervals for the inverse probability weighted estimator, obtained using both methods, achieve coverage rates close to the nominal 95% level. However, slight undercoverage is observed with both approaches, and the extent of undercoverage is comparable between them. In contrast, the sandwich-based confidence intervals for the unweighted estimator perform poorly, particularly for the treatment effect estimate, $\beta_{A}$.
In summary, the asymptotically unbiased nature of the inverse probability weighted estimates, compared to the biased unweighted estimates, further confirms that the simulated data were generated in accordance with the Cox model in (25), and that confounding by $L$ and $W$ was appropriately introduced. R code for implementing this study is provided in Appendix A.
7 Discussion
In this paper, we proposed and demonstrated a principled strategy for parametrically simulating longitudinal data from marginal structural models, leveraging the frugal parameterization (Evans and Didelez, 2024). Our proposed method is highly flexible and can accommodate a variety of practical scenarios, including multiple treatments (for example, primary and drop-in), dynamic treatment regimes, dose-response analyses, multiple outcomes (such as secondary outcomes and competing risks), and censoring. Additionally, it imposes no restrictions on the data-generating process or marginal structural causal models.
In Section 6, we extended this method to survival models and time-to-event outcomes, which are common in clinical trials. Unlike the non-survival case, survival data simulation requires conditioning each time step $k+1$ on surviving through $k$, meaning all parameterized quantities must account for survival. This posed a challenge: while we can specify the distribution of the time-varying variable, say $L_{j}$, conditional on all prior history up to $j-1$, its distribution conditioning on a future survival state at time $k$ ($k>j$) is generally intractable. A very similar issue occurs in the approach proposed by Seaman and Keogh (2024) where Monte Carlo approximations are used to address the problem. While effective in theory, their approach is computationally expensive and imposes restrictions on the types of data that can be generated. We resolved this challenge elegantly by leveraging pair-copulas, so that these conditional quantiles can be iteratively computed through h-functions. Thus, our solution relies solely on evaluating analytic functions, making it highly efficient. This efficiency is crucial for practical applications, such as sample size calculation, which require repeated simulations across a wide range of parameter combinations.
[…]
If enforcing such Markov conditions is absolutely necessary, an alternative pair-copula construction can be adopted. However, such an approach requires computing marginal quantiles of time-varying confounders through numerical integration, which quickly becomes high-dimensional—a well-known computational challenge.
Appendix A Example R code for Section 6.2
Here is an example code to simulate data following the setup in Section 6.2.
1library(data.table)
2library(copula)
3
4sim_func <- function(n = 500,
5 rho1 = -0.5, rho2 = 2,
6 K = 5, betaS = -0.5,
7 betaB2 = -0.5, betaSB2 = 0.3){
8 # copulas
9 copL <- tCopula(param = rho1, df = 5, dim = 2, dispstr = "un")
10 copW <- normalCopula(param = rho2, dim = 2, dispstr = "un")
11 dat <- data.table()
12 dat\$surv <- rep(1,n)
13
14 # baseline vars
15 dat[, B1 := rexp(n, rate = 2)]
16 dat[, B2 := rbinom(n, size = 1, prob = 0.7)]
17 baseline_vars <- c("B1","B2")
18
19 # a data table to store calculated quantiles to be carried over to the next time step
20 comp_dat <- copy(dat[,baseline_vars, with = F])
21
22 for (k in 0:K){
23 n_surv <- sum(dat\$surv, na.rm = T)
24 if (k>0){
25 comp_dat <- comp_dat[Y == 0,]
26 if (k > 1){
27 comp_dat[,grep("_prev\$", colnames(comp_dat), value = TRUE):= NULL]
28 }
29 # reset names
30 setnames(comp_dat,
31 old = setdiff(names(comp_dat), baseline_vars),
32 new = paste0(setdiff(names(comp_dat), baseline_vars), "_prev"))
33 }
34
35 # random quantiles
36 comp_dat[, paste0("F_L",k,"|surv"):= runif(n_surv)]
37 comp_dat[, paste0("F_W",k,"|L",k,"_surv"):= runif(n_surv)]
38 comp_dat[, paste0("F_S",k,"|L",k,"_surv"):= runif(n_surv)]
39 comp_dat[, paste0("F_Y|LW",k,"_surv"):= runif(n_surv)]
40
41 if (k == 0){
42 comp_dat[,L := qgamma(get(paste0("F_L",k,"|surv")), shape = 1 + 0.5*B1 + 0.5*B2, scale = 1)]
43 comp_dat[,W := qbinom(get(paste0("F_W",k,"|L",k,"_surv")), size = 1, prob = expit(-0.2 +0.5 * B1 + 0.5*B2))]
44 } else{
45 comp_dat[,L := qgamma(get(paste0("F_L",k,"|surv")), shape = 1 + 0.5*B1 + 0.5*B2 + 0.1 * L_prev - 0.5 * S_prev, scale = 1)]
46 comp_dat[,W := qbinom(get(paste0("F_W",k,"|L",k,"_surv")), size = 1, prob = expit(-0.2 + 0.5 * B1 + 0.5*B2 + W_prev - 0.6* S_prev))]
47 }
48
49 comp_dat[,S := qbinom(get(paste0("F_S",k,"|L",k,"_surv")),size = 1, prob = expit(-1.5 + 0.5*B1 + 0.5*B2 + 0.5* L + 0.5 * W))]
50
51 if (k >0){
52 #get the distributions of Ls and Ws in the survivors
53 for (j in (1:k)){
54 if (j<k){
55 comp_dat[,paste0("F_W",k-j,"|L",k-j,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_Y|L",k-j,"_W", k-j-1,"_surv_prev")),
56 get(paste0("F_W",k-j,"|L",k-j,"_surv_prev")))),copula = copW)[,2]]
57 comp_dat[,paste0("F_L",k-j,"|surv"):= cCopula(as.matrix(cbind(get(paste0("F_Y|LW",k-j-1,"_surv_prev")),
58 get(paste0("F_L",k-j,"|surv_prev")))),copula = copL)[,2]]
59 } else{
60 comp_dat[,paste0("F_W",k-j,"|L",k-j,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_Y|L",k-j,"_surv_prev")),
61 get(paste0("F_W",k-j,"|L",k-j,"_surv_prev")))),copula = copW)[,2]]
62 comp_dat[,paste0("F_L",k-j,"|surv"):= cCopula(as.matrix(cbind(F_Y_surv_prev,
63 get(paste0("F_L",k-j,"|surv_prev")))),copula = copL)[,2]]
64 }
65 }
66 }
67
68 for (j in (0:k)){
69 if (j<k){
70 comp_dat[,paste0("F_Y|L",k-j,"_W", k-j-1,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_W",k-j,"|L",k-j,"_surv")),
71 get(paste0("F_Y|LW",k-j,"_surv")))),copula = copW,inverse = T)[,2]]
72 comp_dat[,paste0("F_Y|LW",k-j-1,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_L",k-j,"|surv")),
73 get(paste0("F_Y|L",k-j,"_W", k-j-1,"_surv")))),copula = copL,inverse = T)[,2]]
74 }else{
75 comp_dat[,paste0("F_Y|L",k-j,"_surv"):= cCopula(as.matrix(cbind(get(paste0("F_W",k-j,"|L",k-j,"_surv")),
76 get(paste0("F_Y|LW",k-j,"_surv")))),copula = copW,inverse = T)[,2]]
77 comp_dat[,F_Y_surv:= cCopula(as.matrix(cbind(get(paste0("F_L",k-j,"|surv")),
78 get(paste0("F_Y|L",k-j,"_surv")))),copula = copL,inverse = T)[,2]]
79 }
80 }
81
82 # hazard
83 comp_dat[, lambda := exp(-2) *exp(betaS * S + betaB2 * B2 + betaSB2*B2*S)]
84 comp_dat[, T := qexp(F_Y_surv, rate = lambda)]
85 comp_dat[, Y := 1*(T<1)]
86
87 dat[surv == 1, paste0("L","_",k) := comp_dat\$L]
88 dat[surv == 1, paste0("W","_",k) := comp_dat\$W]
89 dat[surv == 1, paste0("S","_",k) := comp_dat\$S]
90 dat[surv == 1, paste0("Y","_",k) := comp_dat\$Y]
91 dat[surv == 1, T := (comp_dat\$T + k)]
92 dat[,surv := 1*(get(paste0("Y_",k)) == 0)]
93 return(dat)
94 }

### v2

@@ Appendix A Extension to Include Time-Varying Latent Variables
We must then obtain $F(y_{k+1}\mid\overline{l}_{j},do(\bar{a}_{k}),y_{k}=0)$ for each $j$ to use in the next iteration. This can be done using (13), and hence, we can move to the $(k+1)$th iteration.
By the principle of induction, the statement is true for all $k\geq 0$. ∎
Appendix C Competing risk
At each time step $k$, after simulating time-varying covariate $L_{k}$ and treatment $A_{k}$, we generate the primary outcome $Y_{k}$ and a competing risk $D_{k}$ using their respective auxiliary variables $\tilde{Y}_{k}$ and $\tilde{D}_{k}$, following the approach outlined in Section 5. For brevity, baseline covariates $\bm{Z}$ continue to be omitted, but in settings where they are present, they can be simulated as the initial step and included in the conditioning set of all subsequent quantities.
Under this simulation framework, we specify:
  • A marginal structural model for the primary event: $F^{*}(\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0)$;
  • A marginal structural model for the competing event: $F^{*}(\tilde{D}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0)$;
  • A bivariate copula parameterized by $\Psi_{k+1}$ to model the dependence between the above interventional margins of $\tilde{Y}_{k+1}$ and $\tilde{D}_{k+1}$;
  • Bivariate copulas that ‘weave’ the multivariate dependence between $Y_{k+1}$ and past covariates $\bar{L}_{k}$, given $do(\bar{A}_{k})$, $Y_{k}=0$ and $D_{k}=0$, parameterized by $\Theta_{k,j}$ for $j=0,\ldots,k$;
  • Bivariate copulas that ‘weave’ the multivariate dependence between $\tilde{D}_{k+1}$ and $\bar{L}_{k}$ conditional on $\tilde{Y}_{k+1}$, $do(\bar{A}_{k})$, $Y_{k}=0$ and $D_{k}=0$, parameterized by $\Phi_{k,j}$ for $j=0,\ldots,k$.
The joint outcome model $p(\tilde{y}_{k+1},\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)$ factorizes as follows:
Primary event model:
$p(\tilde{y}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)$
$=p^{*}(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)\times$
$\prod_{j=0}^{k}c^{*}_{\tilde{Y}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},Y_{k}=0,D_{k}=0}\big{\{}F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0),$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0);\Theta_{k,j}\}.$ (20)
Competing event model:
$p(\tilde{d}_{k+1}$ $\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0)$
$=p(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)\times$
$\prod_{j=0}^{k}c^{*}_{\tilde{D}_{k+1},L_{j}\mid\bar{L}_{j-1},\bar{A}_{k},\tilde{Y}_{k+1}D_{k}=0}\big{\{}F(d_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0),$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0);\Phi{k,j}\}.$ (21)
The first term in (21), the distribution of $\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0$ can be obtained through:
$F(\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},\tilde{y}_{k+1},D_{k}=0)=h\big{\{}F*(\tilde{d}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0),F^{*}(\tilde{y}_{k+1}\,|\,\bar{l}_{k},\bar{a}_{k},Y_{k}=0,D_{k}=0)\big{\}}.$
The challenge lies in computing $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ in (20), and $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ in (21), for $j=0,\ldots,k$. The former can be calculated in the same iterative fashion as described in Section 5. The latter can be derived as:
$F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ $=h\big{\{}F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0),$
$\qquad F(\tilde{y}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0);\Theta_{k,j}\big{\}}.$ (22)
Notably, the two components in the $h$-function in (22) exactly correspond to the those in the primary event model in (20), except with the order of conditioning swapped. Since these values are computed during the simulation of $\tilde{y}_{k+1}$, requiring only $k+1$ additional evaluations of the $h$-functions in (22).
The algorithm for simulating each time step $k$ for an individual can be summarized into:
  1. 1.
    Simulate the time-varying confounder $L_{k}$ and treatment $A_{k}$ from their respective distributions.
  2. 2.
    Compute $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ for $j=0,\dots,k-1$.
  3. 3.
    Compute $F(y_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$ for $j=1,\dots k$, leading to $u_{k+1}=F(\tilde{y}_{k+1}\,|\,z,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$.
  4. 4.
    Compute: $\tilde{y}_{k+1}=F^{*-1}_{\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0,D_{k}=0}(u_{k+1}\,|\,do(\bar{a}_{k}),Y_{k}=0,D_{k}=0)$.
  5. 5.
    Compute $F(l_{j}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$ for $j=0,\ldots,k$ by evaluating $k+1$ $h$-functions following (22).
  6. 6.
    Compute $F(\tilde{d}_{k+1}\,|\,\bar{l}_{j-1},do(\bar{a}_{k}),\tilde{y}_{k+1}=0,D_{k}=0)$ for $j=1,\ldots k$, leading to $v_{k+1}=F(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$.
  7. 7.
    Compute: $\tilde{d}_{k+1}=F_{\tilde{D}_{k+1}\,|\,do(\bar{A}_{k}),\tilde{Y}_{k+1},D_{k}=0}^{*-1}(\tilde{d}_{k+1}\,|\,do(\bar{a}_{k}),\tilde{y}_{k+1},D_{k}=0)$.
This approach requires at most $74k+$ 7 evaluations of analytic functions per individual per time step. An example implementation is available at https://github.com/XiLinStats/LongSiMSM.
Appendix D Additional details of the illustration in Section 5.3
Models for baseline, time-varying covariates and treatment
A baseline:
@@ Appendix C Competing risk
$C\sim\operatorname{Bernoulli}(0.7)$ $W_{0}\sim\operatorname{Bernoulli}(\operatorname{expit}\{-0.2+0.5\,B+0.5\,C\}).$
@@
Parameter Scenarios
Table6 summarizes the parameter values for treatment effect size and confounding scenarios:
@@
Runtime details
Table 7 reports the runtime required to simulate data under the setup for the large causal effect with medium confounding scenario. The simulations were carried out on a local laptop using a single score without parallelization. Generating data for $n$ up to 10,000 units takes only a few seconds, while even for a very large sample size ($n=50,\!000$) and 11 time steps ($K=10$), the runtime remains under one minute. This computational efficiency is crucial for tasks such as method evaluation and sample size calculations in clinical trial design, where simulations must be repeated across various parameter values. For example, determining the sample size might require testing 10 different parameter combinations (such as effect sizes, prevalence of drop-in treatments, or adherence levels) across four sample sizes, each bootstrapped 200 times to estimate statistical power. This requires $10\times 4\times 200=8,\!000$ simulations. Using our method, suppose $n=10,000$ and $K=10$, the total runtime is approximately 13 hours on a single core, which can be reduced to about 2 hours using an 8-core laptop or under 10 minutes on a 50-core CPU through parallelization.
@@
Weight calculation
To fit the Cox marginal model, we employ stabilized inverse probability of treatment weighting, with weights for subject $i$ at time step $k$ calculated as: